OpenStudy (anonymous):

Simplify 2^1/2 x 32^1/2

OpenStudy (anonymous):

Please please help

OpenStudy (anonymous):

Will give medal!!!! I just really need help right now

OpenStudy (whpalmer4):

Is that \[2^{1/2}*32^{1/2}\]If so, you can rewrite it as \[\sqrt{2}*\sqrt{32}\] because \[x^{1/2} = \sqrt{x}\] Next, \[\sqrt{2}*\sqrt{32} = \sqrt{2*32} = \sqrt{2*2*2*2*2*2}\] because 32 = 2*2*2*2*2 and \[\sqrt{a}*\sqrt{b} =\sqrt{a*b}\] Now for every 2*2 under the square root sign, you can place that with a factor of 2 outside the square root sign, and you've got your answer. For example\[\sqrt{3*3*3*3} = 3*3 = 9\]

OpenStudy (anonymous):

Would the answer be 6?

OpenStudy (anonymous):

No wait it would be 8 because you end up getting the square root of 64... Right?